What is the second excitation energy of li+2?
 (1) 12.09ev (2) 48.36ev (3) 108ev (4) 51ev

Grade:12

1 Answers

Brotha
2 years ago
 
The formula for the energy of the electron in n^th orbit is En = -13.6Z^2/n^2 eV.
 
So, for the first orbit (n=1), E1 = -13.6*(3)^2/(1)^2   = -122.4 eV       [Here, Z = 3]
 
For second orbit, n = 2, E2 = -13.6*(3)^2/(3)^2   = -13.6       [Here, Z = 3]
 
The second excitation energy of Li^2+ is the energy when an electron jumps from the first orbit (n = 1) to the third orbit (n = 3), = E2 – E1 = −13.6− (−122.4) =108 eV